OpenStudy (anonymous):

integral of [(2x^2-1)/((4x-1)(x^2+1))dx]

zepdrix (zepdrix):

Is there any particular instructions for solving this? It looks like Partial Fraction Decomposition is required but I just want to check before we try that.

OpenStudy (anonymous):

Yeah, that's what's supposed to be used.

zepdrix (zepdrix):

Hmm so the bottom is factored already, that's nice. Let's see if we can get this broken down correctly.

OpenStudy (anonymous):

I've tried it a few times, but the constants are wrong.

zepdrix (zepdrix):

So here is our initial setup. Understand why the B and C are there? \[\large \frac{2x^2-1}{(4x-1)(x^2+1)} \quad = \quad \frac{A}{4x-1}+\frac{Bx+C}{x^2+1}\]

OpenStudy (anonymous):

Yup, that's what I got

zepdrix (zepdrix):

Multiplying through by that denominator gives us,\[\large 2x^2-1=A(x^2+1)+(Bx+C)(4x-1)\]

zepdrix (zepdrix):

Instead of factoring this all out... let's look for a nice easy X value that might give us some quick information about one of the constants.

OpenStudy (anonymous):

if x=1/4?

zepdrix (zepdrix):

Ok good! :) If we let \(\large x=\frac{1}{4}\), That turns the entire Bx+C term to 0.

OpenStudy (anonymous):

yup, this is where I keep getting the wrong numbers

zepdrix (zepdrix):

\[\large 2\left(\frac{1}{16}\right)-1=A\left(\frac{1}{16}+1\right)\]Ok so far? :o

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

A=-14/17?

zepdrix (zepdrix):

Hmm yah that sounds right. Is that not what it's suppose to be? :o

OpenStudy (anonymous):

-7/34

zepdrix (zepdrix):

We made a boo boo somewhere? Hmm ok gimme sec to think :3

OpenStudy (anonymous):

oh wait..

zepdrix (zepdrix):

It will end up being the constant on top of the 4x-1. Book says -7/34?

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

alright, i forgot about the 1/4 from the integral when i was checking it..

OpenStudy (anonymous):

hold up for a bit?

zepdrix (zepdrix):

k :3

OpenStudy (anonymous):

just checking, does B=48/17 and C=3/17?

zepdrix (zepdrix):

Pshhhh i dunno, i haven't gotten that far yet XD haha

OpenStudy (anonymous):

:O lolk X3

zepdrix (zepdrix):

I got C=3/17 and B=18/51. I should double check my work though... I went through that way too quickly I think.

OpenStudy (anonymous):

Alright, I went through it again but B is still wrong (the final constant when you work through it). :/

OpenStudy (anonymous):

wait how did you get B=18/51?

zepdrix (zepdrix):

I chose x=0 to solve for C. Then having solved for A and C, I plugged those into the problem. From here, I chose x=1 to simplify things down and solved for B.

OpenStudy (anonymous):

now i got B=12/17...

zepdrix (zepdrix):

ok ok gimme sec :3

OpenStudy (anonymous):

I tried it with both -1=-A+C ---> B=12/17 and 2=A+B ---> B=48/17 T^T

zepdrix (zepdrix):

oh you expanded it all out? ew ew :D

OpenStudy (anonymous):

XD

OpenStudy (anonymous):

It's easier for me to see it that way

zepdrix (zepdrix):

\[\large x=1 \qquad \rightarrow \qquad 1=-\frac{14}{17}(2)+\left(B+\frac{3}{17}\right)(3)\]Pshhhh but it plugs in so nicely D":

OpenStudy (anonymous):

O_O It's B=12/17...

zepdrix (zepdrix):

Oh so one of your methods worked? :D ooo exciting!

zepdrix (zepdrix):

Were the A and C correct? :o

zepdrix (zepdrix):

Oh i got 12/17 also, i just didn't calculate correctly the first time.

OpenStudy (anonymous):

Aright, when I expanded it, I got: 2=A+B --->(1) 0=-B+4C --->(2) -1+A-C --->(3) I got: A=-14/17 C=3/17 (Both are correct) When I try to plug A/C into equations (1) or (2) i get different answers. Am I just doing something wrong?

OpenStudy (anonymous):

Yeah, B=12/17 is right., but I'm still curious.

zepdrix (zepdrix):

2=A+4B, I don't think you expanded out those right brackets correctly.

OpenStudy (anonymous):

You're right... I'm sorry, thank you.

zepdrix (zepdrix):

Heh, easy problem to make a mistake on :D sucks.

OpenStudy (anonymous):

I know... -__-

OpenStudy (anonymous):

Thank you so much for your help. I really appreciate it. :)